3(5-a)-4(-1-a) how do i do this problem?
Answer by stanbon(48546)

You have to know the following: a(b+c)=ab+ac
Also, you have to know (negative)*(negative)=positive
So you get 15-3a+1+4a
Now add "like" terms, to get:
a+16
Cheers,
Stan H.